OpenStudy (lgbasallote):

Prove that if m + n and n + p are even integers, where m, n and p are integers, then m + p is even

OpenStudy (lgbasallote):

i suppose first step would be to assume m+n and n+p are even...

OpenStudy (lgbasallote):

so m + n = 2k n + p = 2x

OpenStudy (lgbasallote):

then i suppose i solve for n...

OpenStudy (lgbasallote):

n = 2k - m n = 2x - p

OpenStudy (lgbasallote):

then i suppose i equate them 2k - m = 2x - p

OpenStudy (lgbasallote):

then i suppose this happens 2k - 2x = m - p

OpenStudy (lgbasallote):

then if i factor out... 2 (k - x) = m - p

OpenStudy (lgbasallote):

so... m - p is even but how do i prove m + p is even?

hartnn (hartnn):

m+n = 2x n+p = 2y add them and see whether u can proceed with the proof

OpenStudy (lgbasallote):

ahh yes.. i see it